Please type the answer by computer, do not hand writing the answer

please type the answer by computer, do not hand writing the answer
Answer #1

Assume Cost of Petrol =HK$ P/Litre

Average annual usage=20000Km

Average annual petrol consumption of Corolla=20000/23=869.57 litre

Annual cost of petrol for Corolla=HK $ 869.57*P

Average annual petrol consumption of Prius=20000/27=740.74 litre

Annual cost of petrol for Prius=HK $ 740.74*P

Present Worth Factor (PWF)=(P/A,i,N)=(((1+i)^N)-1)/(i*((1+i)^N))

N=Number of Years=10

i=Interest Rate=5%=0.05

PWF=(P/A,5%,10)=((1.05^10)-1)/(0.05*(1.05^10))=7.721735

Present Value of Annual Cost of Petrol for Corolla=7.721735* HK $ 869.57*P

Present Value of Annual Cost of Petrol for Prius=7.721735* HK $ 740.74*P

Present Worth of Total Cost of Corolla=$(210000+7.721735* 869.57*P)

Present Worth of Total Cost of Prius=$(305000+7.721735* 740.74*P)

At Break even cost of Petrol:

$(210000+7.721735* 869.57*P)= $(305000+7.721735* 740.74*P)

210000+6714.552*P=305000+5719.804*P

P*(6714.552-5719.804)=305000-210000=95000

P*994.748=95000

P=95000/994.748=$95.50

Breakeven Cost of Petrol

95.50

HK $

.(b) Annual Usage =20000X Km

Average annual petrol consumption of Corolla=20000X/23=869.57X litre

Average annual petrol consumption of Prius=20000X/27=740.74X litre

At Break even cost of Petrol:

$(210000+7.721735* 869.57X*P)= $(305000+7.721735* 740.74X*P)

P*X*994.748=95000

P*X=$95.50

Break even Price =P=95.50/X